OpenStudy (anonymous):

Help please! Consider the integral integral of ((x^-2)-4)/x^3 A) Solve this integral by using u-substitution B) Solve the integral another way, using algebra to simplify the integrand first. C) How must your two answers be related? Use algebra to prove this relationship.

hartnn (hartnn):

did you try u = x^-2 -4 ?

hartnn (hartnn):

du = ... ?

OpenStudy (anonymous):

so du = -2x^-3 dx

hartnn (hartnn):

yes du = -2/x^3 dx so, du/(-2) = dx/x^3 plug this in your original integral where you see dx/x^3

OpenStudy (anonymous):

so then that would be integral of u/(-2)du

hartnn (hartnn):

thats correct, now evaluate the integral

OpenStudy (anonymous):

so then it would be... ((x^-2)-4)/-2?

hartnn (hartnn):

naah, integral of u is actually u^2/2 :)

OpenStudy (anonymous):

wait...ummmm...how? :(

hartnn (hartnn):

\(\Large \int x^n dx = \dfrac{x^{n+1}}{n+1}+c\)

OpenStudy (anonymous):

but where does the bottom 2 come from? :(

hartnn (hartnn):

u has the exponent =n =1 denominator = n+1 = 1+1 =2

OpenStudy (anonymous):

oh because -2^1 is still -2

hartnn (hartnn):

yeah

hartnn (hartnn):

but you have u^2 so, \(\Large (u^{-2}-4)^2/2+c\) would be your answer for 1st

hartnn (hartnn):

i mean \(\Large (x^{-2}-4)^2/2+c\)

OpenStudy (anonymous):

how come it's not -2?

hartnn (hartnn):

oops, -2 in the denominator, yes ...

hartnn (hartnn):

\(\Large -(x^{-2}-4)^2/2+c\)

OpenStudy (anonymous):

ok cool so it is integral of u/-2du and becomes u^2/-2 and the du disappears

OpenStudy (anonymous):

leaving -(x^-2-4)/2 + c

OpenStudy (anonymous):

so how would i solve for that algebraically?

hartnn (hartnn):

did i miss the 2 ??? let me write the steps, and ask if you have any doubts in that \(u = x^{-2}-4 \\ du/(-2) = dx/x^3 \\ \int u/(-2)du = (-1/2)\times u^2/2+c = (-u^2/4)+c \\ \Large = [(x^{-2})-4]^2/4+c \)

OpenStudy (anonymous):

wait how does 4 get into the equation? :(

hartnn (hartnn):

the 4 in the denominator ?? -1/2 times u^2/2 = ...

OpenStudy (anonymous):

for the denominator i mean

hartnn (hartnn):

and i missed the negative sign \(u = x^{-2}-4 \\ du/(-2) = dx/x^3 \\ \int u/(-2)du = (-1/2)\times u^2/2+c = (-u^2/4)+c \\ \Large =- [(x^{-2})-4]^2/4+c\)

OpenStudy (anonymous):

i thought the equation for the integral was u/-2du

hartnn (hartnn):

it was

hartnn (hartnn):

for integral of u, we get u^2/2 and that -2 was already in the denominator so, -2 times 2 = -4

OpenStudy (anonymous):

sorry i still don't get where the 2 came from :(

hartnn (hartnn):

the 2 in u^2/2 ??

OpenStudy (anonymous):

so i have this... u= x^(-2)-4 du = -2x^(-3) dx -1/2du=x^(-3)

OpenStudy (anonymous):

yeah :(

OpenStudy (anonymous):

and (-1/2)du=1/(x^3)dx and that becomes... integral of u/(-2)du

OpenStudy (anonymous):

so where does the other 2 come from :(

hartnn (hartnn):

\(\Large \int x^n dx = \dfrac{x^{n+1}}{n+1}+c \\ \Large \int x^1 dx = \dfrac{x^{1+1}}{1+1}+c\)

hartnn (hartnn):

\(\Large (-1/2)\int u^1 dx = (-1/2)\dfrac{u^{1+1}}{1+1}+c = \dfrac{-u^2}{4}+c\)

OpenStudy (anonymous):

oh! i get it now :D

OpenStudy (anonymous):

so how would that be solved algebraically?

hartnn (hartnn):

i am glad you got it :) that was your A) part. now lets do B) part , then we will algebraically simplify it...

hartnn (hartnn):

 right ? can you write each term in the form of x^n ?